# Real-Time Dynamic Robot-Assisted Hand-Object Interaction via Motion Primitives

### v1

@@ V-A Hand Recognition Performance
Figure 6: 3D modeling of the process of converting from the motion "Keep" to the motion "Come". (a) Mesh reconstruction of the hand motion, in which the filled circles represent the joints. (b) Changes of the coordinates of joints during the motion conversion process. Here, "IJ1-4" denotes the four joints of the index finger, respectively.
@@
We first evaluate the effectiveness of the designed 3D hand pose module. Figure 6 demonstrates the 3D modeling of the transition process between two distinct hand motions. In the "Keep" motion, the y-axis coordinates of the four joints remain relatively unchanged. In contrast, during the transition to the "Come" motion, the x-axis coordinates of these joints progressively align. These observations indicate that our 3D hand pose module can accurately capture and model the hand motion changes, providing reliable reference information for the subsequent robot control.
V-B MPM Training
V-B1 Data Preparation
@@
Table I illustrates the pre-defined motion primitives for the proposed ring-wearing task. For each motion, we asked the subject to make and maintain the motion and then collected 2000 consecutive frames of RGB image data, and the subject could make slight motion adjustments during the collection process. Then, these frames were used to extract 3D coordinates of 21 hand joints using the aforementioned pose estimation model. For a single trajectory, we sampled every 10 frames and intercepted a sequence of length 10. As a result, a motion dataset with 1000 samples was obtained, and the date shape of each input is $(10,63)$. Furthermore, we use stratified sampling to divide the data set into a training set (80%) and a validation set (20%) to ensure that each type of action has the same number of samples.
V-B2 Model Training
@@
For the MPM, we leveraged a bidirectional LSTM with 2 recurrent layers, and the input size is (10, 63), and the number of features in the hidden state is 64. The LSTM is followed by a fully connected layer that maps the LSTM output to 5 classes corresponding to different hand motions and a softmax function to output the probability distribution over these classes. We trained the MPM for 100 epochs using a learning rate of 0.00025 and a batch size of 512, and Figure 7 illustrates the model performance changes during training. It can be seen that the model converges quickly and has good generalization performance.
Figure 7: The MPM performance versus training epochs. Accuracy is calculated as the ratio of correctly predicted observations to the total observations, and F1 score is a harmonic mean of precision and recall. The solid line and shaded regions represent the mean and standard deviation computed with five random seeds, respectively.
V-C MPM-based Control and Adaptive Pose Adjustment
V-C1 Motion Primitives Transition
The MPM has been proposed to initiate specific action sets for controlling the robot. Each motion primitive is activated individually. Figure 8 illustrates the motion control using the MPM. In this particular task, the motion of the UR5 along the y-axis is mainly controlled by two motion primitives (supporting Video S1). The robot demonstrates a good dynamic response to the PM commands. In this test scenario, a MP command signal of ’1’ corresponds to the ’Come’ motion, where the robot approaches the human hand, while a MP command of ’2’ signifies a backward motion. The alignment between the detection of the MP command and the y-axis motion showcases effective control of the UR5 robot using the MPM. Note that this example focuses on MPM-based control of motion in one direction. However, the MPM can be expanded to enable similar control over other directional motions, offering versatility in robot control.
Figure 8: The MPM and the control of UR5 arm motion. (a) The PM model is detected by the hand recognition algorithm, and (b) the corresponding motion of the UR5 robot arm in the Y-axis (horizontal) direction.
V-C2 Adaptive Orientation
Figure 9: The adaptive orientation of the UR5 robot arm according to the human input.
Figure 10: The real-world performance of the robot-assisted dressing task in the context of ring-wearing.
@@ V-C2 Adaptive Orientation
In a robot-assisted dressing task, the robot handles environmental and human-induced disturbances. Figure 9 shows the robot’s Z-axis displacement and X-axis angular displacement trajectory. Over 30 seconds, the human bends and then lifts the index finger (see supporting Video S2). During this movement, the robot maintains its grip on the object (ring) and adjusts its TCP orientation in terms of vertical height and pointing angle. The robot achieves a maximum displacement of $\Delta Z$ of 74.23 mm and a rotational angle of 18.0 degrees. These adjustments enable the robot to maintain a suitable position for dressing, validating the effectiveness of the adaptive controller for action alignment discussed in Section IV-C. Note that the robot’s action amplification ratio can be modified to adjust the level of its reactive behavior.
V-D Robot-assisted Ring Wearing
The proposed hand pose estimation and adaptive controller have undergone real-world testing within a robot-assisted dressing task, specifically focusing on a ring-wearing scenario. In this task, the assistive robot is tasked with picking up a ring. Once the object is pinched, the motion of the robot is passively controlled using the MPM, allowing the human to fine-tune the initial positioning for collaboration. Subsequently, triggered by the MPM also, the robot transitions into the dynamic cooperation stage. The robot actively responds to perturbations in hand pose, engaging in collaborative efforts to complete the task of wearing the ring.
Figure 10 illustrates the performance of the robot-assisted ring-wearing task (supporting Video S3). The first and second rows of curves represent the Y-axis and Z-axis motions of the TCP, while the third row depicts the angular displacement around the x-axis, corresponding to the ground coordinate in Figure 5.
@@ V-D Robot-assisted Ring Wearing
Initially, the robot executes a pre-programmed action to pick up the ring. Subsequently, it is controlled by the MPM model, allowing the human operator to guide the end-effector’s approach using the ’Come’ hand pose. Once the operator determines the desired location, they initiate the ’Ring’ motion primitive, transitioning the robot from passive control to active cooperation. From this point (17.6s) onwards, the robot maintains constant movement towards the hand while dynamically compensating for perturbations imposed by the human hand. The dashed blue curve denotes the motion of the CV-based hand pose estimation. As the human adjusts their finger position or location, the robot responds accordingly, ensuring correct ring orientation until successful delivery. At 37.5 seconds, the robot completes the ring-wearing task by releasing the ring, marking the task’s conclusion, and validating the efficiency of the robot in real-world scenarios such as robot-assisted dressing.
VI Conclusion
@@
In this paper, we proposed an adaptive robotic system for robot-assisted hand-object interaction (HOI). We presented a learning-based state estimation that integrates the real-time 3D construction of hand pose. An RNN model has been trained to learn human hand motions. This is used as the motion primitives, which are translated into pre-defined robotic actions in a contactless manner. On the other side, we have designed a closed-loop controller that enables the spatial alignment between the end-effector and the fingering, ensuring that the object is manipulated in the desired way.
@@
In the case study of a ring-wearing task, the adaptive robotic system demonstrates its effectiveness in assisting its human teammate and the robustness against environmental perturbations. Future work can be developed to expand the motion primitive sets and refine interaction dynamics, enhancing the applications of pHRI systems in practical environments.
ACKNOWLEDGMENT
This project is supported by the EU-funded Marie Curie SMART Project (860108).
Ethics Statement
This study involving human participants was conducted in accordance with the principles embodied in the Declaration of Helsinki and local statutory requirements. The research protocol was approved by the Research Ethics Committee of the Department of Engineering, University of Cambridge.
[…]

### v2

@@ 5.1 Hand Recognition Performance
Figure 4: 3D modeling of the process of converting from the motion ”Keep” to the motion ”Come”. (a) Mesh reconstruction of the hand motion, in which the filled circles represent the joints. (b) Changes in the coordinates of joints during the motion conversion process. Here, ”IJ1-4” denotes the four joints of the index finger, respectively.
Figure 5: Real-world performance of the proposed tele-interaction task. (a) Comparison of the timestamps for human hand commands and corresponding robot actions. (b) Adaptive orientation of the UR5 robot arm based on human hand input. (c) Translational and rotational errors between the predicted and actual TCP (Tool Center Point) positions. (d) Multiple stages of the robot-assisted dressing task are demonstrated in the ring-wearing scenario.
@@
We first evaluate the effectiveness of the designed 3D hand pose module. Table 1 illustrates the pre-defined motion primitives while Figure 4 demonstrates the 3D modeling of the transition process between two distinct hand motions. In the ”Keep” motion, the y-axis coordinates of the four joints remain relatively unchanged while under the ”Come” motion the coordinates of these joints progress according to the hand motion. These observations validate the capture and model of the hand motion changes, providing reliable reference information for the subsequent robot control.
5.2 MPM Training
5.2.1 Data Preparation
@@
For each motion, the human participants were instructed to make and maintain the motion and then collected 2000 consecutive frames of RGB image data, and the subject made slight motion adjustments during the collection process. Then, these frames were used to extract 3D coordinates of 21 hand joints using the aforementioned pose estimation model. For a single trajectory, we sampled every 10 frames and intercepted a sequence of length 10. As a result, a motion dataset with 1000 samples was obtained, and the date shape of each input is $(10,63)$. Furthermore, we use stratified sampling to divide the data set into a training set (80%) and a validation set (20%) to ensure that each type of action has the same number of samples.
5.2.2 Model Training
@@
For the MPM, we leveraged a bidirectional LSTM with 2 recurrent layers, and the input size is (10, 63), and the number of features in the hidden state is 64. The LSTM is followed by a fully connected layer that maps the LSTM output to 4 classes corresponding to different hand motions and a softmax function to output the probability distribution over these classes. Finally, we trained the MPM for 100 epochs using a learning rate of 2.5e-4 and a batch size of 512.
5.3 Adaptive Pose Adjustment
5.3.1 Motion Primitives Transition
Each motion primitive is activated individually. Supporting Video S1 demonstrates the motion control using the MPM, where the UR5 robot’s movements along the Y-axis are primarily controlled by two MPs. The robot exhibits a strong dynamic response to these commands, with different MPs triggering actions such as forward and backward movements. Figure 5a illustrates the time delays between the robot’s actions and the human commands during two consecutive come-back MPM commands. It can be seen that at each timestamp, the robot consistently lags behind the human command. For the ’come’ action, the delay is 0.168 seconds, while the ’back’ MPM action experiences a delay of approximately 0.24 seconds. The maximum control delay remains under 0.3 seconds, demonstrating the cobot’s effective command-tracking capabilities within the control architecture.
5.3.2 Adaptive Orientation
@@
The robot handles environmental and human-induced disturbances in real-world implementations. Figure 5b shows the robot’s Z-axis translational and X-axis angular displacement trajectories. Over 30 seconds, the human bends and then lifts the index finger. The robot maintains its grip on the object (ring) and adjusts its TCP orientation in terms of vertical height and pointing angle. The robot achieves a maximum displacement of $\Delta Z$ of 74.23 mm and a rotational angle of 18.0 degrees. These adjustments enable the robot to maintain a suitable position for dressing, validating the effectiveness of the adaptive controller for action alignment discussed in Section 4.3. The error between the predicted TCP and the actual translational TCP is shown in Figure 5c, where the errors fluctuate around zero over time. The maximum absolute translational error is 2.48 mm, while the maximum absolute rotational error reaches 1.37 degrees.
5.4 Robot-assisted Ring Wearing
@@
Figure 5d presents the results of the robot-assisted ring-wearing task (see supporting Video S3). The first and second rows of curves represent the TCP’s motions along the Y and Z axes, respectively, while the third row shows the angular displacement around the x-axis, corresponding to the ground coordinate system shown in Figure 3. In the open-loop hierarchy, the robot first performs a pre-programmed action to pick up the ring. It is then controlled by the MPM model, enabling the human operator to guide the end-effector’s approach using the ’Come’ hand pose. Once the operator specifies the desired location, they activate the ’Ring’ motion primitive, shifting the robot from passive control to active cooperation. From 17.6 seconds onward, the robot operates within the closed-loop hierarchy, continuously adjusting its movement towards the hand and dynamically compensating for any perturbations caused by the human hand. The dashed blue curve represents the motion of the CV-based hand pose estimation. As the human adjusts their finger position or location, the robot adapts in real-time, ensuring the ring remains properly oriented until it is successfully delivered. At 37.5 seconds, the robot completes the ring-wearing task by releasing the ring, thereby demonstrating the effectiveness of the robot in real-world applications such as robot-assisted dressing.
6 CONCLUSION
@@
We proposed a hierarchical procedural framework for robot-assisted hand-object interaction. We presented an open-loop hierarchy that leverages the learning-based state estimation for real-time 3D construction of hand pose. An LSTM-based model has been trained to learn human hand motions. This is used as the motion primitives, which are translated into pre-defined robotic actions in a tele-manipulation manner. Regarding the low-level coordination hierarchy, we have designed a closed-loop controller that enables the spatial alignment between the end-effector and the fingering, ensuring that the object is manipulated collaboratively. Experimental tests have shown that the system exhibits a time delay of $\leq 0.3$ seconds, a translational error of $\leq 2.5$ mm, and a rotational error of $\leq 1.5$ degrees.
@@
In the case study of a ring-wearing task, the adaptive robotic system demonstrates its effectiveness in assisting its human teammate and its robustness against environmental perturbations. Future work can be developed to expand the motion primitive sets and refine interaction dynamics, enhancing the applications of pHRI systems in practical environments.